OpenStudy (anonymous):

Can anyone check my work ? Tell what number you would add to each side of the inequality to solve the inequality f - 6 more than or equal to -3

OpenStudy (anonymous):

f - 6 = -3 -6 +6 --- 9 f = 9

OpenStudy (shamil98):

f - 6 = -3 -3 + 6 = 3 f = 3

OpenStudy (shamil98):

Since it's an inequality it would look like this f - 6 >= -3 f >= 3

OpenStudy (anonymous):

can u help with another ?

OpenStudy (shamil98):

sure.

OpenStudy (anonymous):

a - 3.3 more than or equal to 2.6

OpenStudy (shamil98):

a - 3.3 >= 2.6 a > = 5.9

OpenStudy (anonymous):

ok can u explain it ?

OpenStudy (shamil98):

you add 3.3 to both sides to get a by itself.

OpenStudy (shamil98):

therefore resulting in your answer

OpenStudy (anonymous):

ok how did u know to add 3.3 instead of 2.6

OpenStudy (shamil98):

because 3.3 is on the side with a.

OpenStudy (anonymous):

so whatever is on the same side with the variable gets added

OpenStudy (anonymous):

Ok I have a problem that has the variable by itself

OpenStudy (shamil98):

No, it's easier this way, you could do the question like this a - 3.3 = 2.6 a - 5.9 = 0 -5.9 = -a 5.9 = a

OpenStudy (shamil98):

I just did it that way because its the simplest and easiest way.

OpenStudy (anonymous):

ok how would u do this one 5 > -18 + m

OpenStudy (shamil98):

5 > -18 + m add 18 to both sides 23 > m

OpenStudy (anonymous):

ok why would u add 18

OpenStudy (shamil98):

You want to solve for m. So you want it on one side.

OpenStudy (anonymous):

ok then why not 5 ?

OpenStudy (shamil98):

M - 18 < 5 M and 18 are on the same side , again you could subrtact 5 like i did with an example last problem but its easier this way.

OpenStudy (anonymous):

ok so whatever number is on the same side with the variable is added

OpenStudy (shamil98):

Or subtracted .. m + 5 = 3 -5 -5 m = -2

OpenStudy (anonymous):

ok thanks ! so for v - 4 < -3 I would add 4 to both sides and get v = 1

OpenStudy (shamil98):

Yes.

OpenStudy (anonymous):

OK I think I got it. Thanks ♥

OpenStudy (shamil98):

np.